In a certain orchard there are 4 more rows of trees than there are trees in a row; and if the same number of trees were so arranged that there should be 64 added to each row, the number of rows would be reduced to 4. How many trees are there in the orchard?
Answer by Boreal(15235)

trees in a row=x
rows=x+4
if we had x+64
rows would be 4
The pair is equal
x(x+4)=4x+256
x^2=256
x=16
16 trees in a row, 20 rows, 320 trees ANSWER
add 64 trees to each row, 80 trees in a row, 4 rows, 320 trees. ANSWER